# On the Bounds of Stopping and Cycle Numbers in the Collatz Process

### v15

@@ 2. Collatz Sequence of Stopping
$\begin{array}[]{c|l}m\equiv 3\pmod{4}&\sum_{n_{i}\geq 2}2^{n_{i}}+3\\ \hline\cr 3&3\\ 7&2^{2}+3\\ 11&2^{3}+3\\ 15&2^{3}+2^{2}+3\\ 19&2^{4}+3\\ 23&2^{4}+2^{2}+3\\ 27&2^{4}+2^{3}+3\\ 31&2^{4}+2^{3}+2^{2}+3\\ 35&2^{5}+3\\ 39&2^{5}+2^{2}+3\\ 43&2^{5}+2^{3}+3\\ 47&2^{5}+2^{3}+2^{2}+3\\ 51&2^{5}+2^{4}+3\\ 55&2^{5}+2^{4}+2^{2}+3\\ 59&2^{5}+2^{4}+2^{3}+3\\ 63&2^{5}+2^{4}+2^{3}+2^{2}+3\\ 67&2^{6}+3\\ 71&2^{6}+2^{2}+3\\ 75&2^{6}+2^{3}+3\\ 79&2^{6}+2^{3}+2^{2}+3\\ 83&2^{6}+2^{4}+3\\ 87&2^{6}+2^{4}+2^{2}+3\\ 91&2^{6}+2^{4}+2^{3}+3\\ 95&2^{6}+2^{4}+2^{3}+2^{2}+3\\$ $\vdots$$\vdots\end{array}$
@@
For instance, let m=4i1m=4i-1, then the first two iterates are
@@
The construction in Figure 1 generates all integers $m\equiv 3\pmod{4}$ exactly once. Equivalently, the mapping
@@
Forward ($\Rightarrow$), $q\;\longmapsto\;m,$
We start with $q=11$, and $F_{11}(4i-1)=3^{2}i-1.$
$\Rightarrow$ we have $q=11$..., by rooted sequence tree in Figure 1
$\Rightarrow$ by Collatz process, we have $i=2j$ or $2j+1$, $j=2k$ or $2k+1$, $k=2l$ or $2l+1$, …, $x=2y$ or $2y+1$
$\Rightarrow$ $m=4i-1=2^{s}y+\sum_{n_{i}\geq 2}2^{n_{i}}+3$
$\Rightarrow$ $m=\pmod{4}$
Backward ($\Leftarrow$), $m\;\longmapsto\;q,$
For any, $m\equiv 3\pmod{4}$
$\Rightarrow m=\sum_{n_{i}\geq 2}2^{n_{i}}+3$, by Lemma 2.6
$\Rightarrow$ in m=4i1m=4i-1, based on above $2^{n_{i}}$, we have $i=2j$ or $2j+1$, $j=2k$ or $2k+1$, $k=2l$ or $2l+1$, …,
$\Rightarrow q=1101$..., by $F_{q}(4i-1)$ through inverse Collatz process.
Thus the constructions $q\mapsto m$ and $m\mapsto q$ are inverse to each other, and the correspondence is bijective.

### v16

@@ 2. Collatz Sequence of Stopping
$\begin{array}[]{c|l}4i-1\equiv 3\pmod{4}&\sum_{n_{i}\geq 2}2^{n_{i}}+3\\ \hline\cr 3&3\\ 7&2^{2}+3\\ 11&2^{3}+3\\ 15&2^{3}+2^{2}+3\\ 19&2^{4}+3\\ 23&2^{4}+2^{2}+3\\ 27&2^{4}+2^{3}+3\\ 31&2^{4}+2^{3}+2^{2}+3\\ 35&2^{5}+3\\ 39&2^{5}+2^{2}+3\\ 43&2^{5}+2^{3}+3\\ 47&2^{5}+2^{3}+2^{2}+3\\ 51&2^{5}+2^{4}+3\\ 55&2^{5}+2^{4}+2^{2}+3\\ 59&2^{5}+2^{4}+2^{3}+3\\ 63&2^{5}+2^{4}+2^{3}+2^{2}+3\\ 67&2^{6}+3\\ 71&2^{6}+2^{2}+3\\ 75&2^{6}+2^{3}+3\\ 79&2^{6}+2^{3}+2^{2}+3\\ 83&2^{6}+2^{4}+3\\ 87&2^{6}+2^{4}+2^{2}+3\\ 91&2^{6}+2^{4}+2^{3}+3\\ 95&2^{6}+2^{4}+2^{3}+2^{2}+3\\$ $\vdots$$\vdots\end{array}$
@@
For instance, let $i$ is any integer and $4i-1\equiv 3\pmod{4}$, then the first two iterates are
@@
The construction in Figure 1 generates each integer $m\equiv 3\pmod{4}$ exactly once. Equivalently, the correspondence
@@
We prove the statement by showing that the constructions $q\mapsto m$ and $m\mapsto q$ are well-defined and inverse to each other.
Forward direction ($q\mapsto m$). Starting from the root $q=11$, we have
$F_{11}(4i-1)=3^{2}i-1.$
Any admissible sequence $q=11\cdots$ corresponds to a path in the rooted sequence tree shown in Figure 1. Along this path, the Collatz process, $F_{q}(4i-1)=3^{r}y+F_{q}(m)$, yields successive decompositions
[…]
Consequently, we may write
$4i-1=2^{s}y+\sum_{n_{i}\geq 2}2^{n_{i}}+3=2^{s}y+m,$
which implies
$m\equiv 3\pmod{4}$.
Backward direction ($m\mapsto q$). Let $m$ be any integer satisfying $m\equiv 3\pmod{4}$. By Lemma 2.6, $m$ admits a representation of the form
$m=\sum_{n_{i}\geq 2}2^{n_{i}}+3.$
Writing $4i-1=2^{r}y+m$ and using the powers $2^{n_{i}}$ in the above decomposition, we recover successive relations
$i=2j\text{ or }2j+1,\quad j=2k\text{ or }2k+1,\quad k=2l\text{ or }2l+1,\ \ldots,$
which uniquely determine an admissible sequence
$q=11\cdots$
can be obtained via the inverse Collatz process and the mapping $F_{q}(4i-1)=3^{r}y+F_{q}(m)$.
Since the mappings $q\mapsto m$ and $m\mapsto q$ are inverses of each other, the correspondence between admissible sequences $q$ and integers $m\equiv 3\pmod{4}$ is bijective. ∎
Theorem 2.11.
For any integer $i$, we have
$4i-1\equiv 3\pmod{4},\quad\text{and}\quad F_{q}(4i-1)\neq 4i-1.$
Proof.
From above, we have
$4i-1=2^{s}y+m,$
and
$F_{q}(4i-1)=3^{r}y+F_{q}(m),$
where $y$ is any integer.
Suppose, for the sake of contradiction, that
$4i-1=F_{q}(4i-1).$
Then it follows that
$m=F_{q}(m),$
and hence
$2^{s}y=3^{r}y.$
Since $y\neq 0$, we obtain
$2^{s}=3^{r}.$
This is impossible, because no power of $2$ can equal a power of $3$. Therefore,
$F_{q}(4i-1)\neq 4i-1,$
which completes the proof. ∎
Remark: This implies that for any finite sequence, the existence of a loop is impossible.